# ForgeHLS: A Large-Scale, Open-Source Dataset for High-Level Synthesis

### v2

@@ IV-C Downstream Task 2: Automatic HLS Pragmas Exploration
Our dataset has a large number of C++ kernels without pragma, and corresponding designs with pragma. It is not difficult to envision the potential for utilizing our HLS data to training an automatic pragma inserter based on LLM. Out of 536 ForgeHLS kernels, we use 90% kernels as the training set and reserve the remaining 10% as the test set. We define Pareto designs within our dataset as the dominant solutions in the trade-off between latency and ARU, where no design can improve one of the metrics without sacrificing the other. Then we divide the designs on Pareto curve into three categories: high, medium, and low resource usage. The top one-third of the points in terms of ARU are classified as high resource usage, the middle one-third as medium, and the remaining one-third as low. Lastly, we construct Kernel-Design pairs in the format {"instruction": "optimize for low resource usage and high latency.", "input": %origin code, "output": Pareto design code with HLS pragma consuming low resour%ce usage and high latency"}. By doing so, our trained pragma inserter can generate Pareto-optimal pragma designs for varying resource usage scenarios based on the input kernel code.
@@
We fine-tune the LLaMA3, Qwen, and Mistral models using the LoRA[27] method and evaluate their performance on a subset of 10% kernels from ForgeHLS. For each design generated based on the model’s pragma suggestions, we run the HLS flow and measure the synthesis pass rate. The results, shown in Table IV, indicates that fine-tuning the LLMs on our dataset leads to higher synthesis pass rates for the generated HLS code with inserted pragmas. To further assess the quality of the generated pragma designs, we compare the predicted Pareto designs under three strategies with the ground-truth Pareto designs provided in our dataset. We use the Adjusted Distance to Reference Set (ADRS) as the evaluation metric. A lower ADRS value indicates that the generated designs are closer to the true Pareto front. The ADRS is computed as:
@@
where denotes the total kernel number, denotes the kernel index, denotes the exact Pareto-optimal set of kernel in ForgeHLS, and represents the fine-tuned model predicted Pareto-optimal set in different resource usage strategy for kernel . The term calculates the normalized distance between two design points. refers to the cycle latency of design point, extracted from the report generated by HLS. And refers to ARU of the design point, calculated in equation 1. The expression in ensures that if the proposed method generates a better pragma combination than the ground truth within our dataset, it will be considered to have a zero distance to Pareto curve, which is not penalized. As shown in Table IV, the results demonstrate the effectiveness of fine-tuning LLMs on our dataset for automatic HLS pragma insertion.
@@
Increased HLS Pass Rate. After training, the HLS pass rate increases significantly across all LLMs. For instance, the LLaMA3-7B model trained on ForgeHLS achieves a pass rate of 69%, compared to just 19% for the origin model. The improvement indicates that the trained LLMs are better at inserting valid pragmas. Specifically, model becomes more adept at avoiding invalid pragma insertion locations, improper pragma formats, configuration issues, and C++ syntax errors in the HLS code. This suggests that fine-tuning LLMs on our dataset improves LLMs’ understanding of HLS code format and syntax.
Reduced ADRS. The ADRS is significantly reduced after fine-tuning, indicates that the LLMs, after training, are more adept at generating pragma configurations that are closer to the true Pareto-optimal designs. This further emphasizes the potential of LLMs in DSE, where they can optimize HLS code for better resource usage and latency trade-offs.
@@
Performance on Full ForgeHLS. Despite the improvements in HLS pass rates and ADRS reduction, the ADRS values remain relatively high when training on the full ForgeHLS dataset, compared to only training on synthetic part of ForgeHLS. This highlights the richness and quality of our ForgeHLS, which offers a broader range of HLS code patterns and kernel configurations compared to traditional datasets. The relatively larger ADRS when training on ForgeHLS suggests that more effective learning strategies are needed to fully exploit the diversity and complexity of our dataset. This calls for the development of more advanced models or training methods to leverage the full potential of our dataset for optimizing HLS designs.
@@
V Future Work
Our future work will focus on expanding the ForgeHLS by adding more kernels and incorporating additional DSE. This will further diversify the dataset, enabling more robust models for predicting QoR metrics at scale. There is a clear need for more advanced models that can handle the complexity of such large datasets, and we call for the development of models with better generalization capabilities, such as enhanced GNNs or more powerful LLM architectures. Additionally, exploring new downstream tasks using our dataset, such as automated pragma optimization or hardware-specific design space exploration, will be crucial. Leveraging reinforcement learning or evolutionary algorithms for pragma insertion and resource allocation could help optimize performance further.
VI Conclusion
@@
In this paper, we introduced ForgeHLS, a large-scale, open-source dataset for HLS DSE and QoR prediction. By aggregating diverse real-world and synthetic HLS C++ kernels, ForgeHLS surpasses existing datasets in scale, offering over 536 kernels and 429,276 designs, 10 times larger than prior datasets. Additionally, we developed an automated pragma insertion workflow, enabling direct pragma embedding into the source code to make generated HLS designs eaiser to use. Using this dataset, we conducted two representative downstream tasks: QoR prediction and HLS pragma insertion. These tasks demonstrate the utility of ForgeHLS in HLS optimization and highlight its potential for advancing research in the EDA area.

### v3

@@ 4.2 Downstream Task 2: Automatic HLS Pragmas Exploration
Our dataset has a large number of C kernels without pragma, and corresponding designs with pragma. It is not difficult to envision the potential for utilizing our HLS data to training an automatic pragma inserter based on LLM. For each kernel, we define Pareto designs within our dataset as the dominant solutions in the trade-off between latency and average resource usage (ARU in Eq. 3), where no other designs can improve one of the metrics without sacrificing the other. Then we divide the designs on Pareto curves into three categories: high, medium, and low resource usage. The top one-third of the points in terms of ARU are classified as high resource usage, the middle one-third as medium, and the remaining one-third as low. Lastly, we construct Kernel-Design pairs in a format such as {"instruction": "optimize for low resource usage and high latency.", "input": origin code, "output": Pareto design code with HLS pragma consuming low resource usage and high latency"}.
@@
We fine-tune the LLaMA3, Qwen, and Mistral models on 90% kernels using the LoRA(Hu et al. 2021) method and evaluate their performance on the rest 10% kernels from the same dataset. For each design generated based on the model’s pragma suggestions, we run the HLS flow and measure the synthesis pass rate. To assess the quality of the generated pragma designs, we compare the predicted Pareto designs under three strategies with the ground-truth Pareto designs provided in our dataset. We use the Adjusted Distance to Reference Set (ADRS) as the evaluation metric. A lower ADRS value indicates that the generated designs are closer to the true Pareto front. The ADRS is computed as:
@@
where denotes the total number of kernels, denotes the kernel index, denotes the exact Pareto-optimal set of kernel in ForgeHLS, and represents the Pareto-optimal set predicted by the fine-tuned model for kernel . The term calculates the normalized distance between two design points. refers to the cycle latency of design point, extracted from the report generated by HLS. And refers to ARU of the design point.
@@
Increased HLS Pass Rate. As shown in Table 5, after training, the HLS pass rate increases significantly across all LLMs. For instance, the LLaMA3-7B model trained on ForgeHLS achieves a pass rate of 69%, compared to just 19% for the origin model. The improvement indicates that the trained LLMs are better at inserting valid pragmas. Specifically, the model becomes more adept at avoiding invalid pragma insertion locations, improper pragma formats, configuration issues, and C++ syntax errors in the HLS code. These results indicate that fine-tuned LLMs gain a better understanding of HLS code format and syntax.
Reduced ADRS. Fine-tuning significantly lowers ADRS (e.g., Qwen-7B achieves 38.36%), demonstrating LLMs’ capability to generate pragma configurations closer to Pareto-optimal designs. Critically, these results derive from a one-step exploration paradigm – a simplified DSE setup. The attained low performance at ADRS reveals potential for LLMs in multi-step DSE frameworks, where iterative refinement could further optimize latency-resource trade-offs.
@@
Performance on Full ForgeHLS. The ADRS on the full ForgeHLS is relatively higher than that on the S-ForgeHLS (synthetic part of ForgeHLS). This is because the full dataset contains a broader spectrum of real-world kernels with diverse coding styles and optimization intents, introducing more ambiguity and complexity for learning effective pragma strategies. In contrast, S-ForgeHLS designs exhibit more regular patterns, enabling easier learning and lower ADRS. It also highlights the richness, diversity, and complexity of our ForgeHLS, which offers a broader range of HLS code patterns and kernel configurations. This calls for the development of more advanced models or training methods to leverage the full potential of our dataset for optimizing HLS designs.
@@
5 Conclusion
@@
In this paper, we introduced ForgeHLS, a large-scale, open-source dataset for HLS DSE and QoR prediction. By aggregating diverse real-world and synthetic HLS C++ kernels, ForgeHLS surpasses existing datasets in scale, offering over 846 kernels and 459,850 designs, 10 times larger than prior datasets, consuming over 200k CPU hours during dataset construction. Additionally, we developed an automated pragma insertion workflow, enabling direct pragma embedding into the source code. We conducted two representative downstream tasks: QoR prediction and HLS pragma insertion. These tasks demonstrate the utility of ForgeHLS in HLS optimization and highlight its potential for advancing research in the ML area.
Appendix A Design Space Coverage across Datasets.
As shown in Fig. 4, our dataset demonstrates superior coverage in terms of both latency and resource usage. Our dataset’s larger scale and more comprehensive distribution make it better suited for addressing the diverse challenges encountered in real-world applications. This broader coverage enhances the robustness of model training for tasks like QoR prediction and hardware design optimization.
We evaluate the effectiveness of our approach in expanding the design space coverage, particularly with respect to latency and resource usage, which are essential metrics for model training in downstream tasks such as QoR prediction. For latency, we use the Worst Case Latency extracted from the Vitis HLS report, and for resource usage, we compute the Average Resource Usage (ARU) across four key resource (BRAM, FF, LUT, and DSP). The ARU is calculated using the following equation 4. represents the number of resources for which .
It is also worth noting that HLSyn’s design results in designs that tend to exhibit low resource usage. This leads to a lack of diversity in design solutions. For robust model training, it is crucial to consider not only Pareto-optimal solutions but also non-Pareto solutions. These non-Pareto solutions can also provide valuable insights into the trade-offs between resource usage and latency, contributing to the robustness of the model and its ability to generalize to a broader range of design scenarios.
Appendix B Design Space Coverage of ForgeHLS
We have progressively expanded our dataset to achieve broader kernel coverage. Initially, we used standard HLS benchmark suites, such as CHStone, MachSuite, PolyBench, and Rosetta. We then extended the dataset by incorporating kernels from the Vitis Examples industry database, which significantly improved coverage. Furthermore, the inclusion of kernels generated by LLM Synthetic Code contributed to another substantial increase in design space coverage. As shown in Fig. 5, these kernels, sourced from diverse origins, complement each other in expanding the design space and enriching dataset diversity.
Appendix C Detailed Distribution of ForgeHLS
The detailed distribution of the dataset across different sources is shown in Fig. 6. We present key metrics of our dataset, including the code length, latency, number of pragmas and resource usage. Each data point in these plots represents an individual design, demonstrating the spread of design characteristics across sources. This analysis shows the incremental improvement in coverage as new sources are added, particularly highlighting how LLM-generated synthetic code helps fill gaps in kernel diversity and design complexity. From this distribution, it can be observed that the synthetic code exhibits a relatively rich distribution overall. While the number of pragmas and tokens in synthetic code is slightly lower compared to real-world benchmarks, the feature distribution in terms of resource usage and latency appears comparable. This indicates that its diversity in resource and latency characteristics is sufficient for robust model training.
Appendix D ForgeHLS JSON Format Example
Appendix E Generate Synthetic Data
We input the specifications for the algorithm description and HLS code rules into GPT-4o, which then automatically generate the corresponding code without HLS pragmas. The HLS code rules ensure correct input and output parameter usage, maintain scalability for future pragma additions, enforce strict compliance with the C++11 standard, and prohibit the use of unsupported C/C++ constructs. The unsupported C/C++ constructs include system calls, dynamic memory usage, pointer limitations, recursive functions, undefined behaviors, and virtual functions and pointers.
Appendix F Validate Synthetic Data
We run a single-pass HLS flow to validate the syntactic accuracy of LLM generated code. To validate the functional accuracy, we employ a cross-verification approach leveraging GPT-4o as an assistant. Specifically, we provided GPT-4o with an algorithm description and the corresponding implementation, requesting it to generate a Python version of the same algorithm. This Python implementation served as a reference model for validation. Only when the synthetic code meets both syntactic and functional correctness criteria, we collect it in our ForgeHLS.
Appendix G Full DSE
Following this, we construct a pragma insertion tree to guide the selection of pragma factors. Specifically, we determine whether to enable or disable function inlining and pipelining on the first line inside each function. For the array partition pragma, we choose the target array, the dimension to apply, the partition factor, and the partition type. For loop optimizations, we decide whether to apply loop pipeline, whether to apply unroll, and if unrolling is applied, what unroll factor to choose. It is worth emphasizing that the full DSE does not exhaustively enumerate all pragma permutations. Instead, it follows a set of essential HLS design principles:
If the outer loop in a nested loop structure is annotated with a pipeline pragma, the inner loop will not be assigned an unroll pragma.
The selection of unroll and partition factors is restricted to powers of two, bounded by the loop trip count or array size.
When a branch of the pragma design tree grows excessively large, pruning is applied to retain only configurations with equivalent unroll and partition factors.
By applying these rules, we eliminate invalid pragma combinations, which helps in reducing the dataset size without compromising its quality. A traversal in design tree from the root node to a leaf node represents one pragma configuration. Each path encodes a sequence of pragma decisions, and the resulting leaf node corresponds to a unique design instance. This tree-based representation enables systematic exploration of the design space, where constraints can be enforced dynamically during traversal to avoid infeasible or redundant configurations.
Appendix H Bayesian DSE
To automate the exploration of HLS pragma configurations, we implement a Bayesian optimization framework based on scikit-optimize. Given a C/C++ kernel, the pipeline extracts loop and array structures to construct a discrete design space, where each configuration consists of unroll factors, pipeline toggles, and array partition dimensions.
For each sampled configuration, the system inserts the corresponding HLS pragmas into the source code, synthesizes the design using Vitis HLS, and extracts latency and resource usage from the generated csynth.xml. The resource consumption is normalized with respect to the target FPGA (e.g., xczu9eg-ffvb1156-2-e) across BRAM, LUT, DSP, and FF. The cost function is defined as a balanced trade-off between latency and resource utilization:
where ARU is the average resource usage caculated in equation 4.2
We use Gaussian Process-based optimization with Expected Improvement (EI) as the acquisition function. The search is initialized with 20 random configurations, followed by 40 optimization iterations.
The implementation is submitted as supplementary material and can be invoked via:
Appendix I Experimental Platform Configuration
| FPGA Configuration | |
| Vitis Version | 2023.2 |
| Clock Setting | 10 ns |
| Frequency | 100 MHz |
| FPGA Board | xcu280-fsvh2892-2L-e |
| Available BRAM_18K | 4,032 |
| Available LUT | 1,303,680 |
| Available DSP | 9,024 |
| Available FF | 2,607,360 |
| GPU Task Configuration | |
| GPU Model | NVIDIA A800-SXM4-80GB |
| GPU Quantity | 8 units |
| CPU Task Configuration | |
| Compute Nodes | 3 nodes |
| Cores per Node | 128 cores |
| CPU Model | Intel Xeon Platinum 8462Y+ |
| Total Cores | 384 cores |